OpenStudy (anonymous):

determine if triangle XYZ with coordinates X(2,2) Y(6,7) and Z(7,3) is a right triangle. use evidence to support your claim. if it is not a right triangle what changes could be made to make it a right triangle? medal and fan will be given.

OpenStudy (anonymous):

@zepdrix hi can you help me please

OpenStudy (anonymous):

@ganeshie8 hi can you help me please

OpenStudy (anonymous):

@iambatman can you help me please

OpenStudy (anonymous):

@goformit100 hi can you help me please

OpenStudy (anonymous):

@dumbcow please help me!! can you just tell me what to do please

OpenStudy (anonymous):

@campbell_st can you help me with this please

OpenStudy (anonymous):

@campbell_st do you have an idea how to start bc i was thinking to find the slope or the distance of each line

OpenStudy (dumbcow):

a right triangle would contain a right angle 2 perpendicular lines meet at a right angle find the slope of line between each of the points (3 slopes) and see if any are perpendicular in that they multiply to -1 m1*m2 = -1

OpenStudy (campbell_st):

plot the points using https://www.desmos.com/calculator this will help you find the angle mast likely to be right angled. the find the slope of the lines that make what is possibly a right angle.. lastly show the product of the slopes is -1... which proves the right angle exists

OpenStudy (dumbcow):

another approach is to find the distances between each point and check to see if they model the pythagorean thm a^2 +b^2 = c^2 if they do, then you have a right triangle

OpenStudy (anonymous):

slope formula: m = [y1 - y2] / [x1 - x2] do i use this ? @dumbcow

OpenStudy (campbell_st):

yes... that's as good as any... I just count squares...

OpenStudy (dumbcow):

yes that is slope formula

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

by count squares?

OpenStudy (anonymous):

and who do i give a medal to? bc you both were very helpful

OpenStudy (anonymous):

what do you mean by count squares ? @campbell_st

OpenStudy (campbell_st):

this is what I mean, plot the points and then count the squares to get the slope, rise/run

OpenStudy (campbell_st):

oops the green slope should be YZ not XY

OpenStudy (anonymous):

and how does this prove that it is a right triangle? @campbell_st

OpenStudy (anonymous):

sorry for being annoying i have to explain alot to get full points

OpenStudy (campbell_st):

well you know the slopes of YZ and XZ multiply them together... if the answer is -1 then the segments are perpendicular... any other number and they aren't perpendicular

OpenStudy (anonymous):

yay! thank you!